I need to find the equation for this regression line
x 12 14 10 9 15 8
--------------------
y 29 35 28 23 36 18
Please help
: I need to find the equation for this regression line
x 12 14 10 9 15 8
--------------------
y 29 35 28 23 36 18
Please help

Answer by Fombitz(1756)

Find the average x.
x[ave]=(12+14+10+9+15+8)/6
x[ave]=11.333
Find the average y.
y[ave]=(29+35+28+23+36+18)/6
y[ave]=28.167
For a simple linear regression Y=a+bX, use the following formula to find the slope,
b=(sum( (x-x[ave])(y-y[ave]) ))/(sum((x-x[ave])^2))
b=93.6667/39.333
b=2.381
a=y[ave]-bx[ave]
a=28.167-2.381(11.333)
a=1.178
drawing( 300, 300,-2, 18, -5,50,grid( 1 ), <BR>
circle(12,29,.2),<BR>
circle(14,35,.2),<BR>
circle(10,28,.2),<BR>
circle(9,23,.2),<BR>
circle(15,36,.2),<BR>
circle(8,18,.2),<BR>
blue(line(0,1.178,20,48.80508)) )
Y=2.381X+1.178